OpenStudy (anonymous):

what is the exact value of cos 15

OpenStudy (kymber):

You have to use your calculator, type cos(15)

OpenStudy (shayaan_mustafa):

it can be find out by calculation by using of derivative.

OpenStudy (anonymous):

what is the exact value of cos 15

OpenStudy (kymber):

You may have a scientific calculator at home, or here's a nice online one: http://www.ecalc.com/ type cos(15) into the calculator to get the exact value of cos15.

OpenStudy (anonymous):

i have to break it down

OpenStudy (anonymous):

in steps

OpenStudy (kymber):

Oh. Well I thought you always had to use a calculator. Sorry! Maybe there's a figure to go along with it?

OpenStudy (anonymous):

maybe its ok thank you

OpenStudy (anonymous):

we need the exact value, we can use the identity: \[\cos(A-B) = \cos(A)\cos(B) + \sin(A)\sin(B)\] let A= 45 and B= 30 from an equilateral triangle we can see \[\cos(30) = \sin(60) = \frac{\sqrt{3}}{2}\]and\[\sin(30)=\cos(60) = \frac{1}{2}\] and from this right angled triangle we can see that \[\cos(45) = \sin(45)=\frac{1}{\sqrt{2}}\] put these in the first formula and you can get the exact value